What alkene is the major product formed from each alkyl halide in an El reaction? сну Considering the requirement of an anti- was given the wrong solutions
Answer #1

. CH3 ötte slow Step -Br To ÖH₂ Fast Step (Moderate base) [ RH = 6*H] H - xH = 6*77yote to your posts ochs & Lan [KH = 10 xH)